OpenStudy (anonymous):

Question 1 Simplify the following expression. 2x - 10y + 3x^2 + 7y - 10x. A. 3x2 - 8x - 3y B. 3x2 - 8x - 17y C. 3x2 + 8x - 3y D. 3x2 - 12x + 3y Question 3Evaluate: 32 + (7 - 2) • 4 - 6 over 3 A.24 B. 54 C. 27 D.28

OpenStudy (anonymous):

@kelliegirl33 holy cow... wtf are you writing! a 20 paragraph essay..

OpenStudy (anonymous):

2x - 10y + 3x^2 + 7y - 10x combine like terms : 3x^2 -- there is no other like term with this 2x - 10x = -8x -10y + 7y = -3y now put it all together.. 3x^2 - 8x - 3y ================ 32 + (7 - 2) * 4 - 6/3 32 + 5 * 4 - 6/3 32 + 20 - 2 52 - 2 = 50 ...oops, that is not an answer choice Does this problem have any more parenthesis in it ? sorry it is taking so long...I am kinda stuck on your second question

OpenStudy (anonymous):

No, number 2 is just 3^2+(7-2)x4-6 over three Or 6/3

OpenStudy (anonymous):

I am sorry...I can't seem to get any of your answer choices for the 2nd one

OpenStudy (anonymous):

I can't either..

OpenStudy (austinl):

32 + (7 - 2) • 4 - 6 over 3 \(\displaystyle \frac{32+(7-2)\times4-6}{3}\) \(\displaystyle \frac{32+(5)\times4-6}{3}\) \(\displaystyle \frac{32+20-6}{3}\) \(\displaystyle \frac{46}{3}\) This is what I get.

OpenStudy (anonymous):

No, its just 6 over 3 like 6/3 Not the whole equation

OpenStudy (austinl):

Ahhh, that makes a difference. \(\displaystyle 32+(7−2)×4+\frac{6}{3}\) \(\displaystyle 32+(5)×4+2\) \(\displaystyle 32+20+2\) \(54\)

OpenStudy (anonymous):

I got 46/3 as one of my answers also...that is not an answer choice

OpenStudy (anonymous):

Yeah, like that.. Thank you Austin ! Do you guys mind helping me with a couple more ? This isn't a test btw, its just a review.

OpenStudy (austinl):

sure, why not :)

OpenStudy (anonymous):

wait...it says *4 - 6/3

OpenStudy (anonymous):

oh yeah its - not + ...

OpenStudy (austinl):

Then what the deuce.... there must be an error in the question then.

OpenStudy (anonymous):

I keep getting different answers of which none are the answer choices

OpenStudy (anonymous):

its 4-6/3 Not 4+6/3

OpenStudy (austinl):

It has to be the second one though, it must just be an error. Because if you do that you actually get an answer.

OpenStudy (anonymous):

I kept getting 50 as my answer...lol

OpenStudy (anonymous):

I keep getting 42 ...

OpenStudy (austinl):

According to your given expression, following order of operations (PEMDAS) it should be 50. But I am of the opinion that there is an error in the question. Not by you @lexis98 but more likely the book or site or wherever you got this question.

OpenStudy (anonymous):

I am going to agree with austin...none of my answers are your answer choices

OpenStudy (anonymous):

It's a question on my flvs review practice test. I wrote it just like it gave me.

OpenStudy (anonymous):

Oh well... I'll go with 54!

OpenStudy (anonymous):

Question 6 A truck is rented at $40 per day plus a charge per mile of use. The truck traveled 15 miles in one day and the total charge was $115. Which equation can be used to find the charge x, in dollars, per mile traveled? A. 40x + 15 = 115 B. 40 + 15x = 115 C. 40x + 15x = 115 D. 40 + 15 = 115x Question 11 Which of the following is the most appropriate unit to describe the rate at which data is transferred using the internet? A. Megabits per second, because the independent quantity is time in seconds and dependent quantity is amount of data in megabits. B. Seconds per megabit, because the independent quantity is time in seconds and dependent quantity is amount of data in megabits. C. Seconds per megabit, because the independent quantity is amount of data in megabits and dependent quantity is time in seconds. D. Megabits per second, because the independent quantity is amount of data in megabits and dependent quantity is time in seconds.

OpenStudy (austinl):

Ok, for question 6 we have this basic formula. \(P=dx+40\) Where, P=price d=distance x=price/mile 40 is the flat fee. Do you see anything that fits that?

OpenStudy (anonymous):

These are the answers 40x + 15 = 115 40 + 15x = 115 40x + 15x = 115 40 + 15 = 115x

OpenStudy (austinl):

Is there one that seems to fit my form from the information given?

OpenStudy (anonymous):

40 + 15x = 115 ?

OpenStudy (austinl):

I would agree :)

OpenStudy (anonymous):

Okay, now number 11.

OpenStudy (austinl):

Ok, how would you narrow it down here? You should be able to narrow it down to 2 answer choices.

OpenStudy (anonymous):

A and D?